OpenStudy (anonymous):

For every positive integer, prove that 1+2+...+n= n(n+1)/2

OpenStudy (debbieg):

Inductive proof, maybe?

OpenStudy (asnaseer):

or try reversing the sequence and then adding it to itself, i.e. 1 + 2 + 3 + ... + n n + (n-1) + (n-2) + ... + n -----------------------

OpenStudy (asnaseer):

you may spot a pattern

OpenStudy (asnaseer):

sorry - small mistake: or try reversing the sequence and then adding it to itself, i.e. 1 + 2 + 3 + ... + n n + (n-1) + (n-2) + ... + 1 -----------------------

OpenStudy (debbieg):

if n=2, then n(n+1)/2 = 2(3)/2=3 and 1 + 2 =3 Now assume it hold for k, so the 1+2+....+k=k(k+1)/2 Then show it for k+1, using the result for n=k.

OpenStudy (asnaseer):

@kdekle - you now have two methods to try out - take your pick! :)

OpenStudy (debbieg):

Or better yet... prove it twice! :)

OpenStudy (asnaseer):

he he - too true @DebbieG :)

OpenStudy (anonymous):

so @DebbieG, I should do like Induction to prove that this is correct?